A man of mass 80kg is climbing on a rope whose one end is fixed to the ceiling. If he wishes to go up with an acceleration of 2m/sec ^2, how much force should he apply to the rope?

Grade:11

2 Answers

Abhishek Anand
6 years ago
Answer is 960 N.Solution: As man is climb he exerted a downward force on rope which is tension of the rope.Now in downward direction two forces act ,one is mg of man and other is tension T (say). And in upward direction man moves with acceleration `a` 2m/s^2. As motion is in upward direction so g is -ve, then net force equation is T-mg = maT-80×10=80×2T=800+160T=960 N
Abhishek Anand
6 years ago
Answer is 960 N.Solution: As man is climb he exerted a downward force on rope which is tension of the rope.Now in downward direction two forces act ,one is mg of man and other is tension T (say). And in upward direction man moves with acceleration `a` 2m/s^2. As motion is in upward direction so g is -ve, then net force equation is T-mg = ma ,T-80×10=80×2, T=800+160, T=960 N